Given the following velocities: 380 m/s North, 400 m/s East, 520 m/s 35o North of West, 440 m/s 40o South of West, and 200 m/s 60o South of East. Solve for the resultant velocity.


1
Expert's answer
2021-02-03T16:15:26-0500

Answer

Fitstly we divided into components

So then

X component of velocity

"v_x=400+100-410-337"

=-247m/s

Y component

"v_y=-173+380+298-283"

=222m/s

So resultant velocity

"v=\sqrt{v_x^2+v_y^2}=332.1m/s"